# Starburst-Driven Galactic Outflows

### v1

@@ A.1 Halo model and CR timescales
and velocity dispersion $\sigma=10\,\sigma_{1}\,\text{km s}^{-1}$, where $\rho_{\text{mp}}=\mu\,m_{\text{H}}\,n_{\text{H, mp}}$ is the mid-plane gas density, $\mu=1.4$ is the mean atomic weight, $n_{\text{H, mp}}=n_{0}\,\text{cm}^{-3}$ is the number density of the gas, and where the scale height of the is given by:
@@
where the thermal gas adiabatic index is $\gamma=5/3$, and the CR fluid adiabatic index is $\gamma_{\text{CR}}=4/3$. This configuration creates a layered halo structure, with thermal gas pressure dominating at low altitudes and CR pressure becoming more important at higher altitudes (see the right panel of Fig. 1).
@@ A.2 Outflow model
[…]
@@
We set $\Delta P=0$ to account for the fact that the SBs generally become radiative, and quickly enter a rapidly-cooling wind phase (e.g. Kim & Ostriker 2015; Oku et al. 2022). While the expansion of such rapidly cooling winds is still formally energy-driven, the coupling between the hot interior and the shell leads to dynamics that are equivalent to those of a momentum-driven wind, but with a slightly boosted momentum injection rate in comparison to that at the source (Lancaster et al. 2024). We account for this boost by means of the energy efficiency factor $\varepsilon_{\text{w}}$.
We model the CRs as a non-thermal fluid with an adiabatic index of $\gamma_{\text{CR}}=4/3$, assuming a uniform pressure distribution immediately behind the shock. The CR pressure then evolves as:
$P_{\text{CR, in}}=3\left(\gamma_{\text{CR}}-1\right)\frac{f_{\text{CR}}\varepsilon_{w}\left(\dot{E}_{\text{SB}}/4\pi\right)t}{z^{3}}\ .$ (13)
@@
$v_{\infty}=\frac{v_{\text{ej}}}{2\sqrt{1-f_{\text{CR}}}}\left(\delta+\sqrt{\delta^{2}+2f_{\text{CR}}}\right),$ (16)
@@
$v_{\infty}^{\text{weak}}\rightarrow 66\,f_{\text{CR}}\varepsilon_{\text{w, -2}}\,E_{51}\,\mathcal{R}_{-3}\,\sigma_{1}^{-4}\,\text{km s}^{-1}\ .$ (19)
However, in the strong source limit where $\mathcal{R}_{\text{SN}}\gg\mathcal{R}_{\text{c}}$, the outflow velocities are much higher, approaching:
$v_{\infty}^{\text{strong}}\rightarrow 10^{3}\,\varepsilon_{w,-2}^{1/2}\,E_{51}^{1/2}\,M_{\text{ej, 0}}^{-1/2}\frac{\sqrt{1+f_{\text{CR}}}+\sqrt{1-f_{\text{CR}}}}{2}\,\text{km s}^{-1}\ .$ (20)
Appendix B Model Limitations and Assumptions
@@
Density Profile and Gravitational Field. The density profile in eq. (7) represents an isothermal, single-component atmosphere in vertical hydrostatic equilibrium, providing a reasonably accurate description near the mid-plane. However, the presence of molecular gas and a young stellar disk could create a deeper potential well, leading to a more compact density profile that may affect the early stages of shock breakout. Moreover, explicitly modeling the gravitational potential of different galaxy components (e.g. the bulge, disk and dark-matter halo) could modify flow properties (see Shimoda & Inutsuka 2022), and this can be affected by the detailed gas distribution throughout the galaxy and inner halo. However, the outflow properties at high altitudes (above $\sim 100$ kpc) are unlikely to be significantly impacted if the overall gas surface density and wind loading remain unchanged.
@@
Halo Gas Properties. In the halo, the multi-phase CGM contributes a diffuse gas background that exerts additional (ram) pressure, which can counteract outflow expansion (Shin et al. 2021). While the CGM is diffuse, it is expected to follow a shallow density profile that would slow the outflow’s expansion once the swept-up CGM mass becomes comparable to $M_{\infty}$. This occurs at a height of $z_{\text{CGM}}\sim 5\,\sigma_{1}\,n_{0}^{-1/6}\,n_{\text{CGM, -3}}^{-1/3}\,\text{kpc}$, where $n_{\text{H, CGM}}=10^{-3}\,n_{\text{CGM, -3}}\,\text{cm}^{-3}$ is the number density of hydrogen in the CGM.
@@
Ejecta Mass. For the asymptotic steady-state solution, we consider a limit where the outflow has traveled sufficiently far above the disk, yet is not old enough for its mass to be dominated by the ejecta. The timescale for the outflow to become ejecta-dominated is $t_{\text{ED}}\sim 4\pi M_{\infty}/\dot{M}_{\text{SB}}\sim 12.6\,\sigma_{1}^{3}\,n_{0}^{-1/2}\,M_{\text{ej, 0}}^{-1}\,\mathcal{R}_{-3}^{-1}\,\text{Gyr}$, which is considerably longer than the lifetime of the starburst driving the outflow. We therefore consider our results to be robust against this approximation.
@@
Thin-Shell and Sector Approximation. While the thin-shell approximation is well-suited for radiative blastwaves or those propagating through media with positive density gradients, it becomes increasingly crude in environments with steep negative density gradients. This is because the mass, energy, and momentum distributions behind the shock broaden significantly (see e.g. Laumbach & Probstein 1969; Koo & McKee 1990). Despite the steep density gradient considered in this work, our model focuses on a radiative, momentum-driven wind, where the thin-shell approximation is expected to remain reasonably accurate. The sector approximation, on the other hand, is most reliable when the local shock surface remains relatively flat. This approximation can break down if adjacent streamlines begin to diverge significantly, e.g. due to sudden deflections. However, since observed outflows generally exhibit large opening angles, this approximation is likely to have only a negligible impact on our findings.
@@
Treatment of CRs. Our treatment of CRs involves a number of simplifications. While we do not explicitly account for CR cooling or interactions within the flow, this is well justified given the relevant timescales (see Appendix A.1). However, future studies with more detailed CR propagation modeling may yield different quantitative results, particularly if a detailed CR transport model within the halo is included. Such models could alter the distribution of halo CRs (Recchia et al. 2021), and reveal the microphysical impacts of CRs on outflows, including re-acceleration processes and interactions with complex magnetic field structures at the outflow–halo interface (see, e.g. Hopkins et al. 2023). Additionally, our study does not consider the spectral evolution of CR particles, which could influence the coupling between CRs and the outflowing wind fluid, potentially modifying their driving efficiency and altering observable CR emission signatures. Future work incorporating these effects could provide a clearer understanding of how CRs shape the development of outflows in CR-rich galaxy halos.

### v2

@@ A.1 Halo model and CR timescales
and velocity dispersion $\sigma=10\,\sigma_{1}\,\text{km s}^{-1}$, where $\rho_{\text{mp}}=\mu\,m_{\text{H}}\,n_{\text{H, mp}}$ is the mid-plane gas density, $\mu=1.4$ is the mean atomic weight, $n_{\text{H, mp}}=n_{0}\,\text{cm}^{-3}$ is the number density of the gas, and where the disk scale-height is given by:
@@
where $f_{\text{CR}}$ is the fraction of energy supplied by the starburst to CRs at the galactic mid-plane, $\gamma=5/3$ is the thermal gas adiabatic index, and $\gamma=4/3$ is the CR fluid adiabatic index. This configuration creates a layered halo structure, with thermal gas pressure dominating at low altitudes and CR pressure becoming more important at higher altitudes (see the right panel of Fig. 1).
@@ A.2 Outflow model
Here $z=z\left(t\right)$ is the shock radius at time $t$ after the onset of shock expansion, $\dot{z}$ and $\ddot{z}$ are the instantaneous expansion velocity and acceleration of a thin shell following the expanding shock, $M=\int\rho z^{2}dz$ is the swept-up mass, and $\Delta P$ and $\Delta P_{\text{CR}}$ are the thermal and non-thermal pressure differences between the shocked and the un-shocked gas, respectively. The gravitational acceleration, $g<0$, for a single-component, isothermal slab in vertical hydrostatic equilibrium is given by:
@@
In the sector approximation (Laumbach & Probstein 1969), we follow the one-zone dynamics of a shell-segment per unit angle, i.e. along a single streamline, which captures its dilution in three dimensions due to the growth of the surface area of the shell-segment as the shock expands.
We set $\Delta P=0$ to account for the fact that the SBs generally become radiative, and quickly enter a rapidly-cooling wind phase (e.g. Kim & Ostriker 2015; Oku et al. 2022). While the expansion of such rapidly cooling winds is still formally energy-driven, the coupling between the hot interior and the shell leads to dynamics that are equivalent to those of a momentum-driven wind, but with a slightly boosted momentum injection rate in comparison to that at the source (Lancaster et al. 2024). We account for this boost, as well as losses, by means of the energy efficiency factor $\varepsilon_{\text{w}}$.
We model CRs as a non-thermal fluid with an adiabatic index of $\gamma_{\text{CR}}=4/3$, which is advected with the outflowing gas (Girichidis et al. 2024), assuming a uniform pressure distribution immediately behind the shock. The CR pressure then evolves as:
$P_{\text{CR, in}}=3\left(\gamma_{\text{CR}}-1\right)\frac{f_{\text{CR}}\varepsilon_{\text{w, CR}}\left(\dot{E}_{\text{SB}}/4\pi\right)t}{z^{3}}\ ,$ (13)
where $\varepsilon_{\text{w, CR}}=\zeta_{\rm str,CR}\;\!\zeta_{\rm cal,CR}\sim 0.01$ is an efficiency factor for the CR energy. This parametrizes the combined effects of energy losses due to CR streaming, $\zeta_{\rm str,CR}\sim 0.1$ (Chan et al. 2019; Hopkins et al. 2022), which accounts for their enhancement by stronger magnetic fields near the galactic disk (Buck et al. 2020), and CR attenuation by hadronic collisions, characterized by $\zeta_{\rm cal,CR}$.
Galactic starbursts are often CR calorimeters, meaning that a substantial amount of the CR energy is lost to hadronic pp collisions within their ISM (Krumholz et al. 2020; Werhahn et al. 2023). The CR calorimetric fraction has been reported to vary between 0.8 and 0.99 in nearby starbursts (Krumholz et al. 2020). However, such strong calorimetry within the outflow itself would generally only arise before the outflow has cleared a path through the ISM of its host, i.e. the ¡10Myr break-out timescale (see Appendix B) compared with the ¿100 Myr duration of a wind-sustaining starburst (e.g. Di Matteo et al. 2008). A fiducial choice of $\zeta_{\rm cal,CR}\sim 0.1$ is therefore conservative.
@@
$v_{\infty}=\frac{v_{\text{ej}}}{2\sqrt{1-f_{\text{CR}}}}\left(\delta+\sqrt{\delta^{2}+2\,\left(\varepsilon_{\text{w, CR}}/\varepsilon_{\text{w}}\right)\,f_{\text{CR}}}\right),$ (16)
@@
$v_{\infty}^{\text{weak}}\rightarrow 66\,f_{\text{CR}}\,\varepsilon_{\text{w, CR, -2}}\,E_{51}\,\mathcal{R}_{-3}\,\sigma_{1}^{-4}\,\text{km s}^{-1}\ ,$ (19)
where $\varepsilon_{w,{\rm CR},-2}\equiv\varepsilon_{w,{\rm CR}}/10^{-2}$ and ${\mathcal{R}_{-3}}\equiv{\mathcal{R}_{\rm SN}}\,{\rm kyr}^{-1}.$
However, in the strong source limit where $\mathcal{R}_{\text{SN}}\gg\mathcal{R}_{\text{c}}$ and $\delta\simeq\sqrt{1-f_{\rm CR}}$, the outflow velocities are much higher, approaching:
$v_{\infty}^{\text{strong}}$ $\rightarrow$ $5\times 10^{2}\,\varepsilon_{w,-2}^{1/2}\,E_{51}^{1/2}\,M_{\text{ej, 0}}^{-1/2}\left(\sqrt{1-f_{\text{CR}}}\right$. (20)
$+\left.\sqrt{1+\left(2\,\varepsilon_{\text{w, CR}}/\varepsilon_{\text{w}}-1\right)f_{\text{CR}}}\right)\,\text{km s}^{-1}\ .$
Appendix B Timescales
To assess the validity of the underlying assumptions of our model, we check that the timescales related to the shock break-out from the ISM and the time it takes to reach the steady state solution are short enough to have a negligible effect on the overall dynamics.
The time it takes for a continuously driven SB to become radiative and dissipate the majority of its thermal energy is on the order of the shell-formation timescale (Oku et al. 2022)
$t_{\text{sf}}\sim 0.12\,\mathcal{R}_{-3}^{0.28}\,n_{0}^{-0.71}\,\text{Myr}\leavevmode\nobreak\ ,$ (21)
which is usually much shorter than the corresponding time it would take for an adiabatic blastwave to reach the galactic scale height and break out; on the order of
$t_{\text{s}}^{\text{adiab.}}\sim 1.7\,\sigma_{1}^{5/3}\,n_{0}^{-1/2}\,\mathcal{R}_{\text{SN, }-3}^{-1/3}\,\text{Myr}\leavevmode\nobreak\ .$ (22)
Thus, unless they are extremely powerful, the dynamics of starburst-driven SBs prior to shock break-out can be described by a momentum-driven wind solution $z\propto t^{1/2}$ and the time it takes for the shock to break out is longer due to the radiative losses; on the order of
$t_{\text{s}}^{\text{rad.}}\sim 9.3\,\sigma_{1}^{2}\,n_{0}^{-1/2}\,\left(1-f_{\text{CR}}\right)^{-1/4}\,M_{\text{ej, }0}^{-1/4}\,\mathcal{R}_{\text{SN, }-3}^{-1/2}\,\text{Myr}\leavevmode\nobreak\ .$ (23)
Prior to shock break-out the dynamics of the shock are well described by the expressions for momentum-driven winds in uniform media (Oku et al. 2022; Lancaster et al. 2024) in which the losses are explicitly accounted for. We can thus set $\varepsilon_{\text{w}}=1$ for these pre-break-out expressions. At the time of shock break-out the shock speed will have dropped dramatically to only about
$v_{\text{s}}^{\text{break-out}}\sim 18\,\sigma_{1}^{-1}\,\left(1-f_{\text{CR}}\right)^{1/4}\,M_{\text{ej, }0}^{1/4}\,\mathcal{R}_{-3}^{1/2}\,\text{km s}^{-1}\leavevmode\nobreak\ ,$ (24)
and the shock is re-accelerated before it reaches its asymptotic speed. In the case of a weak CR-driven wind the acceleration after break-out is approximately $\propto v_{\text{s}}^{-1}$, leading to rapid acceleration up to the asymptotic speed Eq. 19 within only
$t_{\text{accel}}^{\text{weak}}\sim\frac{v_{\text{s}}}{\dot{v}_{\text{s}}}\sim 15\,\sigma_{1}^{-1/2}\,n_{\text{0}}^{-1/2}\,\varepsilon_{\text{w, CR, }-2}^{-1/2}\,f_{\text{CR}}^{-1/2}\,M_{\text{ej, }0}^{3/8}\,\mathcal{R}_{-3}^{1/4}\,\text{Myr}\leavevmode\nobreak\ ,$ (25)
where we took the geometric average of the acceleration between $v_{\text{s}}^{\text{break-out}}$ and $v_{\infty}^{\text{weak}}$. In this time frame the outflow can already reach a few kpc above the disk. In the strong-source limit the acceleration after break-out is $\propto\left(v_{\infty}^{\text{strong}}-v_{\text{s}}\right)$, which leads to a moderate, but still relatively fast acceleration to the asymptotic speed Eq. 20, within
$t_{\text{accel}}^{\text{strong}}\sim 224\,\sigma_{1}^{2}\,n_{0}^{-1/2}\,\varepsilon_{\text{w, }-2}\,\left(1-f_{\text{CR}}\right)^{-1/4}\,M_{\text{ej, }0}^{-1/4}\,\mathcal{R}_{\text{SN, }-3}^{-1/2}\,\text{Myr}\leavevmode\nobreak\ .$ (26)
In this time frame the wind may have already expanded many 10s of kpcs into the halo, while stronger winds may reach their asymptotic speed closer to the galaxy.
These timescales may be compared to the typical duration of a starburst episode on the order of $t_{\text{starburst}}\sim 10\,\text{Myr\,}$(Rathjen et al. 2023), which suggests that starburst driven winds are established over the course of multiple short bursts, each lasting only a fraction of the lifetime of the wind. The SN rate $\mathcal{R_{\text{SN}}}$ in our model thus represents the average over a series of distinct starbursts.
Appendix C Model Limitations and Assumptions
@@
The density profile in eq. (7) represents an isothermal, single-component atmosphere in vertical hydrostatic equilibrium, providing a reasonably accurate description near the mid-plane. However, the presence of molecular gas and a young stellar disk could create a deeper potential well, leading to a more compact density profile that may affect the early stages of shock breakout. Moreover, explicitly modeling the gravitational potential of different galaxy components (e.g. the bulge, disk and dark-matter halo) could modify flow properties (see Shimoda & Inutsuka 2022), and this can be affected by the detailed gas distribution throughout the galaxy and inner halo. However, the outflow properties at high altitudes (above $\sim 100$ kpc) are unlikely to be significantly impacted if the overall gas surface density and wind loading remain unchanged.
Halo Gas Properties.
@@
In the halo, the multi-phase CGM contributes a diffuse gas background that exerts additional (ram) pressure, which can counteract outflow expansion (Shin et al. 2021). While the CGM is diffuse, it is expected to follow a shallow density profile that would slow the outflow’s expansion once the swept-up CGM mass becomes comparable to $M_{\infty}$. This occurs at a height of $z_{\text{CGM}}\sim 5\,\sigma_{1}\,n_{0}^{-1/6}\,n_{\text{CGM, -3}}^{-1/3}\,\text{kpc}$, where $n_{\text{H, CGM}}=10^{-3}\,n_{\text{CGM, -3}}\,\text{cm}^{-3}$ is the number density of hydrogen in the CGM.
Ejecta Mass.
@@
For the asymptotic steady-state solution, we consider a limit where the outflow has traveled sufficiently far above the disk, yet is not old enough for its mass to be dominated by the ejecta. The timescale for the outflow to become ejecta-dominated is $t_{\text{ED}}\sim 4\pi M_{\infty}/\dot{M}_{\text{SB}}\sim 12.6\,\sigma_{1}^{3}\,n_{0}^{-1/2}\,M_{\text{ej, 0}}^{-1}\,\mathcal{R}_{-3}^{-1}\,\text{Gyr}$, which is considerably longer than the lifetime of the starburst driving the outflow. We therefore consider our results to be robust against this approximation.
Thin-Shell and Sector Approximation.
@@
While the thin-shell approximation is well-suited for radiative blastwaves or those propagating through media with positive density gradients, it becomes increasingly crude in environments with steep negative density gradients. This is because the mass, energy, and momentum distributions behind the shock broaden significantly (see e.g. Laumbach & Probstein 1969; Koo & McKee 1990). Despite the steep density gradient considered in this work, our model focuses on a radiative, momentum-driven wind, where the thin-shell approximation is expected to remain reasonably accurate. The sector approximation, on the other hand, is most reliable when the local shock surface remains relatively flat. This approximation can break down if adjacent streamlines begin to diverge significantly, e.g. due to sudden deflections. However, since observed outflows generally exhibit large opening angles, this approximation is likely to have only a negligible impact on our findings.
Treatment of CRs.
@@
Our treatment of CRs involves a number of simplifications. While we do not explicitly account for CR cooling or interactions within the flow, this is well justified given the relevant timescales (see Appendix A.1), at least in the context of low mass galaxies. In more massive galaxies with denser gas, it becomes increasingly difficult for shocks driven by stellar feedback to break-out of the disk, where cooling losses do dominate, rendering CR-driven galactic outflows ineffective in such systems with halo masses above $\gtrsim 10^{12}\,\text{M}_{\odot}$ (Jacob et al. 2018; Girichidis et al. 2024). However, future studies with more detailed CR propagation modeling may yield different quantitative results, particularly if a detailed CR transport model within the halo is included. Such models could alter the distribution of halo CRs (Recchia et al. 2021), and reveal the microphysical impacts of CRs on outflows, including re-acceleration processes and interactions with complex magnetic field structures at the outflow–halo interface (see, e.g. Hopkins et al. 2023). Additionally, our study does not consider the spectral evolution of CR particles, which could influence the coupling between CRs and the outflowing wind fluid, potentially modifying their driving efficiency and altering observable CR emission signatures. Future work incorporating these effects could provide a clearer understanding of how CRs shape the development of outflows in CR-rich galaxy halos.